# On Arratia’s coupling and the Dirichlet law for the factors of a random integer

### v2

@@ 1.1. Application to the distribution of factorizations of random integers
When $u\in[0,1/2]$, we have $(1+u_{i}\log x)^{1-\alpha_{i}}(1+(1-u_{i})\log x)^{\alpha_{i}}\geqslant(0.5\log x)^{\alpha_{i}}$. Similarly, when $u\in[1/2,1]$, we have $(1+u_{i}\log x)^{1-\alpha_{i}}(1+(1-u_{i})\log x)^{\alpha_{i}}\geqslant(0.5\log x)^{1-\alpha_{i}}$. We thus find that the expression in the big-Oh in Theorem 2 is $\leqslant(0.5\log x)^{-\min\{\alpha_{1},\dots,\alpha_{k},1-\alpha_{1},\dots,1-\alpha_{k}\}}$ uniformly in $\mathbf{u}\in[0,1]^{k-1}$. Since $1-\alpha_{j}\geqslant\alpha_{i}$ for all $i\neq j$ by our assumption that $\alpha_{1}+\cdots+\alpha_{k}=1$, we conclude that the error term in Theorem 2 is $\ll_{\bm{\alpha}}(\log x)^{-\min\{\alpha_{1},\dots,\alpha_{k}\}}$, thus recovering Leung’s estimate (1.7) when $f$ lies in the class $\mathcal{F}_{k}(\bm{\alpha})$.
The proof of Theorem 2 is based on a 1987 result of Donnelly and Tavaré [8], who proved the following probabilistic version of Leung’s theorem: If $\mathbf{V}=(V_{1},V_{2},\ldots)$ is a Poisson–Dirichlet process and $(C_{i})_{i\geqslant 1}$ is a sequence of i.i.d. random variables supported on $\{1,\ldots,k\}$ satisfying $\mathbb{P}[C_{i}=j]=\alpha_{j}$, then
(1.8) $\Bigg{(}\sum_{i:\ C_{i}=1}V_{i},\ldots,\sum_{i:\ C_{i}=k}V_{i}\Bigg{)}$
follows exactly the Dirichlet distribution $\operatorname{Dir}(\bm{\alpha})$. In 1998, Arratia [1] used this result to show that the left-hand side of (1.6) is $\frac{2}{\pi}\arcsin\sqrt{u}+o(1)$ as $x\to\infty$ with probabilistic methods. We use the coupling to bridge between the distribution of (1.8) and Theorem 2 and get an explicit error term.
Remarks.
Here is a brief heuristic about the shape of the error term we obtain in Theorem 2. Let $\bm{\delta}_{x}\coloneqq(\frac{\log D_{x,1}}{\log N_{x}},\ldots,\frac{\log D_{x,k}}{\log N_{x}})$. There exists a coupling between $\bm{\delta}_{x}$ and the random vector (1.8) such that their distance is of typical size $\asymp\frac{1}{\log x}$. For each $j$, the marginal distribution of the $j$th component of $\operatorname{Dir}(\bm{\alpha})$ is $\mathrm{Beta}(\alpha_{j},1-\alpha_{j})$, which is why we get the error term of Theorem 2 when none of the $u_{i}$’s are close to either 00 or $1$.
When one of the $u_{i}$’s is close to 00 or $1$, the behavior changes completely in the error term of Theorem 2, because the vector $\bm{\delta}_{x}$ has a discrete distribution (compared to the continuous distribution $\operatorname{Dir}(\bm{\alpha})$). For instance, if $u_{i}<\log 2/\log x$ for some $i$, then the relation $D_{x,i}\leqslant N_{x}^{u_{i}}$ implies that $D_{x,i}=1$.
@@ 1.2. Structure of the paper
  • In Section 9, we use the coupling of Section 2 to construct a coupling of the random $k$-factorization $\mathbf{D}_{x}$ and an analogous $k$-partition of the Poisson–Dirichlet distribution. We then use this coupling to reduce Theorem 2 to estimating two boundary events, one involving number-theoretic objects and the other one using pure probabilistic objects.
@@
The readers interested only in Theorem 2 need to only understand Section 2 from Part I.
@@ 1.3. Notation
Moreover, for $n\in\mathbb{N}$, we let $s(n)$ denote its largest square-full divisor. Also, we let $n^{\flat}\coloneqq n/s(n)$ and we note that $n^{\flat}$ is square-free and co-prime to $s(n)$.
@@
If $P$ is some proposition, then the indicator function $\mathbbm{1}_{P}$ will be equal to 1 if $P$ is true and 00 if $P$ is false. For a set or an event $A$, we will sometimes write $\mathbbm{1}_{A}(\omega)$ to mean $\mathbbm{1}_{\omega\in A}$.
@@ 2. The coupling
(2.1) $\lambda_{j}=\log q_{j}+O(1/(\log q_{j})^{2}).$
Indeed, Mertens’s theorems and the Prime Number Theorem [12, Theorems 3.4 and 8.1] yield
$\sum_{p^{k}\leqslant y}\frac{1}{kp^{k}}=\log_{2}y+\gamma+O(1/(\log y)^{3})\qquad(y\geqslant 2).$
A detailed proof of this estimate is given in Proposition A.2. Taking $y=q_{j}$ in it proves (2.1).
@@
    Construct the sequence of random prime powers or ones $(Q_{i})_{i\geqslant 1}$ by letting $Q_{i}\coloneqq e^{h(L_{i}\log x)}$. Note that $Q_{i}=1$ whenever $L_{i}\leqslant\frac{e^{-\gamma}}{\log x}$ so there are only finitely many $Q_{i}$’s that are prime powers.
@@
    • $\mathbb{P}\big{[}M_{x}\neq N_{x}\big{]}=d_{\mathrm{TV}}(\mu_{x},\nu_{x})$
    with $d_{\mathrm{TV}}$ being the total variation distance defined in (B.1).
@@ 2.1. Reducing Theorem 1 to three lemmas and a proposition
$\Theta_{x}\coloneqq\sum_{i\geqslant 1}r(V_{i}\log x).$
@@
Indeed, if we let $a=n^{\flat}$ and $b=s(n)$, then
@@
where we used our assumption that $\alpha<1$. Since we have assumed that $\beta<1/2$, the sum over $b$ converges, thus completing the proof. ∎
@@
Let $S\coloneqq\sum_{i\geqslant 1}|\log P_{i}-V_{i}\log x|$. We always have the trivial bound $S\leqslant 2\log x$. This bound and Lemma 2.1 gives us
@@ 3. The $\ell^{1}$ distance within the coupling
For two non-increasing sequences $(x_{i})_{i\geqslant 1}$ and $(y_{i})_{i\geqslant 1}$ of real numbers, and for any two permutations $\sigma,\rho\colon\mathbb{N}\to\mathbb{N}$, we have
@@
Recall that the definitions of the sequence $(Q_{i})_{i\geqslant 1}$, the extra prime $P_{\mathtt{extra}}$ and $M_{x}$. We create another sequence of primes or ones $(\widetilde{P}_{i})_{i\geqslant 1}$ in the following way:
@@ 4. Another realization of the GEM distribution
    We denote by $\mathscr{R}$ the Poisson process on $\mathbb{R}_{>0}^{2}$ that has intensity measure $e^{-wy}\,\mathrm{d}w\,\mathrm{d}y$. Without loss of generality, we may assume that the $w$-coordinates of the points of $\mathscr{R}$ are always distinct.
@@
Applying the map $T_{x}(s)\coloneqq\log_{2}x-\log s$ to the points of $\{S_{i}\,:\,i\in\mathbb{Z}\}$ gives a homogeneous Poisson process on the real line with constant rate $1$, by using the Mapping Theorem (Proposition B.3) with Lemma 4.2. Furthermore, we have $T_{x}(S_{0})<0\leqslant T_{x}(S_{1})$ and $(T_{x}(S_{i}))_{i\in\mathbb{Z}}$ increasing. Therefore, $T_{x}(S_{1})$ and $T_{x}(S_{i+1})-T_{x}(S_{i})$ for all $i\geqslant 1$ are independent exponential random variables of parameter $1$. If $X$ is a standard exponential random variable, then $1-e^{-X}$ is a uniform random variable in $[0,1]$. We conclude that $1-\frac{S_{1}}{\log x},\ \frac{Y_{1}}{S_{1}},\ \frac{Y_{2}}{S_{2}},\,\ldots$ are independent uniform random variables in $[0,1]$. The proposition follows by the characterization of the GEM distribution described in the introduction. ∎
@@ 5. An integer-friendly version of $J_{x}$
This random integer $J^{*}_{x}$ was used by Arratia in [2], and he estimated its distribution in his Lemma 2. For the sake of completeness, we repeat his proof in Lemma 5.2 below. But first, we need the following preliminary result on a particular sum of independent Poisson random variables.
Lemma 5.1.
Let $\lambda\in(0,1)$, let $(X_{k})_{k\geqslant 1}$ be a sequence of independent Poisson random variables such that $\mathbb{E}[X_{k}]=\frac{\lambda^{k}}{k}$, and let $Z\coloneqq\sum_{k\geqslant 1}k\cdot X_{k}$. Then we have
$\mathbb{P}\big{[}Z=\ell\big{]}=(1-\lambda)\cdot\lambda^{\ell}\qquad\text{for}\ \ell=0,1,2\dots$
Proof using generating functions.
Let $s\in\mathbb{C}$ with $|s|<\lambda^{-1}$. We have that $\mathbb{E}[s^{X_{k}}]=\exp(\frac{\lambda^{k}}{k}\cdot(s-1))$. With the independence of the $X_{k}$’s, we have
$\mathbb{E}[s^{Z}]=\prod_{k=1}^{\infty}\exp(\tfrac{\lambda^{k}}{k}\cdot(s^{k}-1))=\exp\big{(}-\log(1-\lambda s)+\log(1-\lambda)\big{)}=\sum_{j=0}^{\infty}(1-\lambda)\lambda^{j}\cdot s^{j}.$
We finally recover the probability mass function of $Z$ using
$\mathbb{P}\big{[}Z=\ell\big{]}=\frac{1}{2\pi i}\oint_{|s|=1}\frac{\mathbb{E}[s^{Z}]}{s^{\ell+1}}\,\mathrm{d}s.\qed$
Proof using combinatorics.
For the case $\ell\geqslant 1$, we have
$\mathbb{P}\big{[}Z=\ell\big{]}$ $=\sum_{\begin{subarray}{c}m_{1},\ldots,m_{\ell}\geqslant 0\\ \sum jm_{j}=\ell\end{subarray}}\Bigg{(}\prod_{j=1}^{\ell}\mathbb{P}[X_{j}=m_{j}]\prod_{j=\ell+1}^{\infty}\mathbb{P}[X_{j}=0]\Bigg{)}$
$=\exp\!\bigg{(}-\sum_{j=1}^{\infty}\frac{\lambda^{j}}{j}\bigg{)}\cdot\lambda^{\ell}\cdot\sum_{\begin{subarray}{c}m_{1},\ldots,m_{\ell}\geqslant 0\\ \sum jm_{j}=\ell\end{subarray}}\prod_{j=1}^{\ell}\frac{(1/j)^{m_{j}}}{m_{j}!}.$
By a classical formula due to Cauchy [16, Proposition 1.3.2], the expression $\ell!\cdot\prod_{j=1}^{\ell}\frac{(1/j)^{m_{j}}}{m_{j}!}$ equals the number of permutations in $S_{\ell}$ with $m_{j}$ cycles of length $j$ for all $j$. In particular,
$\sum_{\begin{subarray}{c}m_{1},\ldots,m_{\ell}\geqslant 0\\ \sum jm_{j}=\ell\end{subarray}}\prod_{j=1}^{\ell}\frac{(1/j)^{m_{j}}}{m_{j}!}=1,$
whence $\mathbb{P}[Z=\ell]=(1-\lambda)\cdot\lambda^{\ell}$, as claimed. Finally, for the case $\ell=0$, we recover $\mathbb{P}[Z=0]$ using the formula $\mathbb{P}[Z=0]=1-\mathbb{P}[Z\geqslant 1]$. ∎
@@
Let $t>0$ and let $q$ be a prime power, and consider the random variable $N_{q}(t)$ that counts the number of points $(T^{*}_{i},Q^{*}_{i})$ with $T^{*}_{i}>t$ and $Q^{*}_{i}=q$. Moreover, let $I^{*}_{t}\coloneqq\prod q^{N_{q}(t)}$ with the product being over all prime powers $q$, and let $Z_{p}(t)\coloneqq\sum_{k=1}^{\infty}kN_{p^{k}}(t)$. Thus, we have
$I^{*}_{t}=\prod_{p}p^{Z_{p}(t)}.$
The family of random variables $(Z_{p}(t))_{p}$ is independent since $\mathscr{R}^{*}$ is a Poisson process. We also know that each random variable $N_{p^{k}}(t)$ follows a Poisson distribution with parameter $\frac{1}{kp^{k(1+t)}}$. Hence, $\mathbb{P}\big{[}Z_{p}(t)=\ell\big{]}=(1-p^{-1-t})p^{-\ell(1+t)}$ by Lemma 5.1 for all $\ell\geqslant 0$, and we have
$\mathbb{P}\big{[}I^{*}_{t}=j\big{]}=\prod_{p}\mathbb{P}\big{[}Z_{p}(t)=\nu_{p}(j)\big{]}=\frac{j^{-1-t}}{\zeta(1+t)}$
for all $j\geqslant 1$. All points of $\mathscr{R}^{*}$ have distinct $T^{*}$-coordinates with probability one. In this situation, we have that $J_{x}^{*}=j$ if, and only if, there exists exactly one point $(T^{*},Q^{*})\in\mathscr{R}^{*}$ such that $I^{*}_{T^{*}}=j$ and $Q^{*}>\frac{x}{j}$. Thus, we have
@@
almost surely. Taking expectations on both sides and using the Mecke equation (Proposition B.6), we get the distribution of $J_{x}^{*}$:
@@
We want to estimate this integral. Let $S(u)\coloneqq\sum_{q\leqslant u}\frac{\Lambda(q)}{q}$. We have $S(u)=\log u+O(1)$ for $u\geqslant 1$ by Mertens’ estimate [12, Theorem 3.4(a)]. We use partial summation to get
@@ 6. When $J_{x}$ and $J_{x}^{*}$ are different
  • $\mathcal{E}_{1}=\big{\{}S_{1}<\log x-R_{x}-r_{0},\ S_{0}>\log x+R_{x}+2r_{0}\big{\}}$, where $R_{x}\coloneqq\sum_{i\geqslant 2}r(Y_{i})$.
@@
Recall that $J_{x}=\prod_{i\geqslant 1}e^{h(Y_{j})}$ and assume that $\mathcal{E}_{1}$ occurs. Then
@@
Recall that we constructed $N_{x}$ in the coupling such that $\mathbb{P}[M_{x}\neq N_{x}]=d_{\mathrm{TV}}(\mu_{x},\nu_{x})$ with $\mu_{x}$ and $\nu_{x}$ being the distribution of $M_{x}$ and $N_{x}$ respectively. Note that $M_{x}>x$ must imply that $J_{x}>x$, which means that $\mathcal{E}_{1}$ cannot happen by (6.1). Hence, using Lemma 6.3, we see that it is enough to show that
@@
  • $\mathcal{B}_{1}=\{J_{x}=J_{x}^{*}\leqslant x/2\}$;
  • $\mathcal{B}_{2}=\{M_{x}\in A\}$;
  • $\mathcal{B}_{3}=\{\mbox{$J_{x}^{*}=j$ and $U^{\prime}_{1}\in A_{j}$ for some $j\leqslant x/2$}\}$.
@@
$\Big{|}\mathbb{P}[\mathcal{B}_{2}]-\mathbb{P}[\mathcal{B}_{3}]\Big{|}\leqslant\mathbb{P}[\mathcal{B}_{1}^{c}]\leqslant\mathbb{P}\big{[}J_{x}\neq J_{x}^{*}\big{]}+\mathbb{P}\big{[}x/2<J_{x}^{*}\leqslant x\big{]}\ll\frac{1}{\log x}.$
@@
Since $\theta(t)/t=1+O\big{(}(\log t)^{-2}\big{)}$ for all $t\geqslant 2$ by the Prime Number Theorem [12, Theorem 8.1], we have

### v3

@@ 1.1. Application to the distribution of factorizations of random integers
Here is a brief heuristic about the shape of the error term we obtain in Theorem 2. Suppose that $N_{x}\neq 1$ and let $\bm{\delta}_{f,x}\coloneqq(\frac{\log D_{f,x,1}}{\log N_{x}},\ldots,\frac{\log D_{f,x,k}}{\log N_{x}})$. There exists a coupling between $\bm{\delta}_{f,x}$ and the random vector (1.9) such that their distance is of typical size $\asymp\frac{1}{\log x}$. For each $j$, the marginal distribution of the $j$th component of $\operatorname{Dir}(\bm{\alpha})$ is $\mathrm{Beta}(\alpha_{j},1-\alpha_{j})$, which is why we get the error term of Theorem 2.
@@ 1.2. Structure of the paper
  • In Section 9, we use the coupling of Section 2 to construct a coupling of the random $k$-factorization $\mathbf{D}_{f,x}$ and an analogous $k$-partition of the components of the Poisson–Dirichlet process. We then use this coupling to reduce Theorem 2 to estimating two boundary events, one involving number-theoretic objects and the other one using pure probabilistic objects.
@@
The readers interested only in Theorem 2 do not need to go beyond Section 2 in Part I.
@@ 1.3. Notation
Moreover, for $n\in\mathbb{N}$, we let $s(n)$ denote its largest square-full divisor. Also, we let $n^{\flat}\coloneqq n/s(n)$ and we note that $n^{\flat}$ is always square-free and co-prime to $s(n)$.
@@
If $P$ is some proposition, then the indicator function $\mathbbm{1}_{P}$ will be equal to 1 if $P$ is true and 00 if $P$ is false.
@@ 2. The coupling
(2.1) $\lambda_{j}=\log q_{j}+O(1/(\log q_{j})^{2})$
by a stronger version of Mertens’s estimate (Proposition A.2).
@@
    Construct the sequence of random prime powers or ones $(Q_{i})_{i\geqslant 1}$ by letting $Q_{i}\coloneqq e^{h(L_{i}\log x)}$. Note that we have $Q_{i}=1$ whenever $L_{i}\leqslant\frac{e^{-\gamma}}{\log x}$ so there are only finitely many $Q_{i}$’s that are prime powers.
@@
    • $\mathbb{P}\big{[}M_{x}\neq N_{x}\big{]}=\mathrm{d}_{\mathrm{TV}}(\mu_{x},\nu_{x})$
    with $\mathrm{d}_{\mathrm{TV}}$ being the total variation distance defined in (B.1).
@@ 2.1. Reducing Theorem 1 to three lemmas and a proposition
(2.5) $\Theta_{x}\coloneqq\sum_{i\geqslant 1}r(V_{i}\log x).$
@@
Indeed, if we let $b=s(n)$ and $a=n^{\flat}=n/b$, then
@@
where we used our assumption that $\alpha<1$. Since we have assumed that $\beta<1/2$, the last sum over $b$ converges, thus completing the proof. ∎
@@
Let $S\coloneqq\sum_{i\geqslant 1}|\log P_{i}-V_{i}\log x|$. The proof that $\mathbb{E}[S]\gg 1$ is explained in the remark after the statement of the theorem in the introduction. For the upper bound, we first note that we always have the trivial bound $S\leqslant 2\log x$. This bound and Lemma 2.1 gives us
@@ 3. The $\ell^{1}$ distance within the coupling
For any two non-increasing sequences $(x_{i})_{i\geqslant 1}$ and $(y_{i})_{i\geqslant 1}$ of real numbers, and for any two permutations $\sigma,\rho\colon\mathbb{N}\to\mathbb{N}$, we have
@@
Recall the definitions of the sequence of prime powers $(Q_{i})_{i\geqslant 1}$, the extra prime $P_{\mathtt{extra}}$ and $M_{x}$. We create another sequence of primes or ones $(\widetilde{P}_{i})_{i\geqslant 1}$ in the following way:
@@ 4. Another realization of the GEM distribution
    We denote by $\mathscr{R}$ the Poisson process on $\mathbb{R}_{>0}^{2}$ that has intensity measure $e^{-wy}\,\mathrm{d}w\,\mathrm{d}y$. We may assume that the $w$-coordinates of the points of $\mathscr{R}$ are all distinct since this happens almost surely.
@@
Applying the map $T_{x}(s)\coloneqq\log_{2}x-\log s$ to the points of $\{S_{i}\,:\,i\in\mathbb{Z}\}$ yields a homogeneous Poisson process on the real line with constant rate $1$, by the Mapping Theorem (Proposition B.3) and Lemma 4.2. Furthermore, we have $T_{x}(S_{0})<0\leqslant T_{x}(S_{1})$ and $(T_{x}(S_{i}))_{i\in\mathbb{Z}}$ increasing. Therefore, $T_{x}(S_{1})$ and $T_{x}(S_{i+1})-T_{x}(S_{i})$ for all $i\geqslant 1$ are independent exponential random variables of parameter $1$. If $X$ is a standard exponential random variable, then $1-e^{-X}$ is a uniform random variable in $[0,1]$. We conclude that $1-\frac{S_{1}}{\log x},\ \frac{Y_{1}}{S_{1}},\ \frac{Y_{2}}{S_{2}},\,\ldots$ are independent uniform random variables in $[0,1]$. The proposition follows by the characterization of the GEM distribution described in the introduction. ∎
@@ 5. An integer-friendly version of $J_{x}$
This random integer $J^{*}_{x}$ was used by Arratia in [2], and he estimated its distribution in his Lemma 2. For the sake of completeness, we give a proof for this estimation.
@@
Let $t>0$ and let $q$ be a prime power, and consider the random variable $N_{q}(t)$ that counts the number of points $(T^{*}_{i},Q^{*}_{i})$ with $T^{*}_{i}>t$ and $Q^{*}_{i}=q$. Note that $N_{q}(t)$ has Poisson distribution with parameter $\frac{\Lambda(q)}{q^{1+t}\log q}$. Moreover, let $I^{*}_{t}\coloneqq\prod q^{N_{q}(t)}$ with the product being over all prime powers $q$. On one hand, when $t$ is fixed, the expectation $\mathbb{E}[(I^{*}_{t})^{-s}]$ can be expanded as a Dirichlet series whose $i$th coefficient equals $\mathbb{P}[I^{*}_{t}=i]$. This Dirichlet series converges absolutely for $\text{Re}(s)\geqslant 0$. On the other hand, we have
$\mathbb{E}[(I^{*}_{t})^{-s}]=\prod_{q}\mathbb{E}\big{[}q^{-sN_{q}(t)}\big{]}=\prod_{q}\exp\!\Big{(}\tfrac{\Lambda(q)}{q^{1+t}\log q}(q^{-s}-1)\Big{)}=\frac{\zeta(1+t+s)}{\zeta(1+t)}.$
It follows that $\mathbb{P}[I^{*}_{t}=i]=\frac{i^{-1-t}}{\zeta(1+t)}$ for all $i\geqslant 1$. All points of $\mathscr{R}^{*}$ have distinct $T^{*}$-coordinates with probability one. In this situation, we have that $J_{x}^{*}=j$ if, and only if, there exists exactly one point $(T^{*},Q^{*})\in\mathscr{R}^{*}$ such that $I^{*}_{T^{*}}=j$ and $Q^{*}>\frac{x}{j}$. Thus, we have
@@
almost surely. Taking expectations on both sides and using the Mecke equation (Proposition B.6), we obtain the distribution of $J_{x}^{*}$:
@@
We want to estimate this integral. Let $S(u)\coloneqq\sum_{q\leqslant u}\frac{\Lambda(q)}{q}$. We have $S(u)=\log u+O(1)$ for $u\geqslant 1$ by Mertens’s estimate [14, Theorem 3.4(a)]. We use partial summation to get
@@ 6. When $J_{x}$ and $J_{x}^{*}$ are different
  • $\mathcal{E}_{1}\coloneqq\big{\{}S_{1}<\log x-R_{x}-r_{0},\ S_{0}>\log x+R_{x}+2r_{0}\big{\}}$, where $R_{x}\coloneqq\sum_{i\geqslant 2}r(Y_{i})$.
@@
Recall that $J_{x}=\prod_{i\geqslant 1}e^{h(Y_{i})}$ and assume that $\mathcal{E}_{1}$ occurs. Then
@@
Recall that we constructed $N_{x}$ in the coupling such that $\mathbb{P}[M_{x}\neq N_{x}]=\mathrm{d}_{\mathrm{TV}}(\mu_{x},\nu_{x})$ with $\mu_{x}$ and $\nu_{x}$ being the distribution of $M_{x}$ and $N_{x}$ respectively. Note that $M_{x}>x$ must imply that $J_{x}>x$, which means that $\mathcal{E}_{1}$ cannot happen in that situation by (6.1). Hence, using Lemma 6.3, we see that it is enough to show that
@@
  • $\mathcal{B}_{1}\coloneqq\{J_{x}=J_{x}^{*}\leqslant x/2\}$;
  • $\mathcal{B}_{2}\coloneqq\{M_{x}\in A\}$;
  • $\mathcal{B}_{3}\coloneqq\{\mbox{$J_{x}^{*}=j$ and $U^{\prime}_{1}\in A_{j}$ for some $j\leqslant x/2$}\}$.
@@
$\Big{|}\mathbb{P}[\mathcal{B}_{2}]-\mathbb{P}[\mathcal{B}_{3}]\Big{|}\leqslant\mathbb{P}[\mathcal{B}_{1}^{c}]\leqslant\mathbb{P}\big{[}J_{x}\neq J_{x}^{*}\big{]}+\mathbb{P}\big{[}x/2<J_{x}^{*}\leqslant x\big{]}\ll\frac{1}{\log x}$
@@
Since $\theta(t)/t=1+O\big{(}(\log t)^{-2}\big{)}$ for all $t\geqslant 2$ by the Prime Number Theorem [14, Theorem 8.1], we have